Supply: q = (1/4)p

Demand: q = 6 − (1/2)p

What is the Deadweight Loss under a price floor of $10?

(A) $2.
(B) $3.
(C) $6.
(D) None of the above.
Answer #1

Demand Supply : : 0 = 6 - P/2 Og = P/4 At equlibruem , 6- P = P14 6= 3p14 p=8 , 0=2 Oy tot 74 1=1/4 Р- Ч 17 Week 07 SUNDAY WE

Thus, correct option: (b) $3